Joan Gunderson rented the same car model twice from Hertz, which rents this car model for a daily fee plus an addition charge per mile driven. Joan recalls that the car rented for 5 days and driven for 300 miles cost her $178, while the same model car rented for 4 days and driven for 500 miles cost $197. Find the daily fee and find the mileage charge.
Answer by richwmiller(17219)

y=mx+b
b= daily fee
m =mileage fee
178=m*300+5*b,
197=m*500+4*b
daily fee = $23, mileage fee = $.21
check
178=.21*300+5*23,
178=63+115
197=.21*500+4*23
197=105+92
ok